A glass worker needs to prepare 20.54 L diluted solution of pH 2.10 but only has 25.88 ml concentrated bottle of acid . HF is a very toxic and deadly acid, but great for cleaning glass. On the label it is written that 25.11 g of HF was used to prepare the concentrated acid. What volume of concentrated acid is required to make the diluted solution?
1
Expert's answer
2020-12-01T09:11:51-0500

pH of the solution = 2.10

So, –log [H+] = 2.10

Or, [H+] = 7.943×10^–3

Considering complete dissociation of HF, we can write, [HF] = [H+]

So, [HF] = 7.943×10^–3 M

So, molarity of diluted HF solution

= 7.943×10^–3 M


Volume of acid solution, V = 25.88 ml

Mass of HF , W = 25.11 g

Molar mass of HF, M = 20 g/mol

No. Of moles of HF = (25.11/20) mol

= 1.2555 mol

25.88 ml solution contain 1.2555 mol HF

1 ml solution contain (1.2555/25.88) mol HF

1000 ml solution contain

= (1.2555/25.88)×1000 mol HF

= 48.51 mole

So, molarity of the concentrated HF solution

= 48.51 M


We know, V1×S1=V2×S2

Or, V2 = V1×S1/S2


V1 & V2 are the volume of diluted and concentrated solution respectively

S1 & S2 are the volume of diluted and concentrated solution respectively


So, V2 = (20.54×7.943×10^–3)/48.51

= 3.37×10^–3 mL


Hence, 3.37×10^–3 mL of concentrated acid is required to make the diluted solution.